Computer Graphics and a New Gibbs Phenomenon for Fourier–Bessel Series

Alfred Gray Mark A. Pinsky

Abstract

We report the existence of a Gibbs-like phenomenon at points of continuity in the expansion of functions in Fourier–Bessel series.